OpenStudy (anonymous):

Applied Calculus - At a production level of 350 silver ski pairs, a manufacturing company estimates a marginal cost of $60 per ski pair and an average cost of $85 per ski pair. A) Estimate the average cost if 357 ski pairs are produced. B) Is the average cost per ski pair increasing, decreasing, or staying the same at a production level of 350 ski pairs? Provide reasoning.

OpenStudy (dumbcow):

Because the marginal cost is less than average cost when level is at 350, by producing more skis the avg cost will decrease since the cost of making those extra skis is cheaper. cost of producing the 7 extra skis is 60*7 = 420 total cost = 350*85 +420 = 30,170 new avg cost = 30,170/357 = 84.5